OpenStudy (anonymous):

Find x: 25(15^x-1) = (3^x-1)

OpenStudy (anonymous):

\[25(15^{x-1}) = 3^{x-1}\]

OpenStudy (anonymous):

Could someone explain what I should do, step by step?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

Yes, I know that the answer is x = -1. But, how?

OpenStudy (anonymous):

is it correct..?

OpenStudy (anonymous):

Yes! :)

OpenStudy (anonymous):

first consider x-1=y

OpenStudy (anonymous):

then it becomes 25(15^y) = (3^y)

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

15^y=(3^y)(5^y)

OpenStudy (anonymous):

cancelling 3^y on both sides...

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

we find 25 (5^y)=1

OpenStudy (anonymous):

thus y=-2=x-1

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

hence proved... :)